OpenStudy (anonymous):

Determine whether the sequence is arithmetic or geometric. Sequence 1: –10, 20, – 40, 80, ... Sequence 2: 15, – 5, – 25, – 45, ... Which of the following statements are true regarding Sequence 1 and Sequence 2. Both sequences are arithmetic. Sequence 1 is arithmetic and Sequence 2 is geometric. Both sequences are geometric. Sequence 2 is arithmetic and Sequence 1 is geometric.

OpenStudy (anonymous):

What is the difference between arithmetic and geometric?

OpenStudy (anonymous):

Arithmetic has a constant difference between terms and geometric has a constant ratio between terms @zimmah

OpenStudy (radar):

Which sequence has a constant difference between terms?

OpenStudy (anonymous):

I think it would be pretty hard to prove which one of them has a constant difference and which one doesnt. Given the options you have for an answer, that's pretty much all you need to do.

OpenStudy (anonymous):

Sequence 1 @radar

OpenStudy (anonymous):

Pretty hard = pretty easy, lol. I was merging two sentences in a wrong way. It's not very hard or it's pretty easy, not pretty hard :p

OpenStudy (anonymous):

-10 -x = 20 Can you solve x here?

OpenStudy (radar):

Lets see -10 - 20 = -30 -20 -40 = - 60 -40 - 80 =-120 does that difference look constant?

OpenStudy (anonymous):

yeah it's going up by 20 @radar

OpenStudy (radar):

-30 to for the first pair difference -60 for the second pair differenc -120 for the next pair difference That does not look like a constant 20 to me......??

OpenStudy (radar):

You still think the first sequence is arithmetic sequence. Look at the evidence

OpenStudy (anonymous):

30

OpenStudy (radar):

What you must look at and answer this question, is the difference between each pair is the same.

OpenStudy (radar):

is 20, 30, and 60 the same ??

OpenStudy (radar):

If they are not the same, then it is not an arithmetic sequence.

OpenStudy (radar):

Does 20 = 30? does 30 = 60? Don't think so. So they must be different. The difference is not constant.